# Large language models can be zero-shot anomaly detectors for time series?

### v1

@@ 4 SigLLM: Detecting Anomalies in Signals using Large Language Models
4.1 Prompter: Finding Anomalies through Prompt Engineering
@@
Under this prompt, the LLM generates a list of values it delineates as point-wise anomalies. It is noteworthy that the GPT-3.5-turbo model is capable of directly outputting anomalous indices using the prompt presented in Table 1, while Mistral lacks this ability as shown in trial #5. To maintain consistency across our experiments, we conducted experiments on both models using the same prompt mentioned above.
As explained in Section III.A, we adopt the rolling windows method, segmenting the time series into rolling windows before inputting it into the LLMs. For each window, we generate 10 samples from the output probability distribution. For each sample containing values deemed anomalous by LLMs, we collect all indices of the window corresponding to those values. Then, the 10 lists of indices are merged together: if an index appears in at least $\alpha$ percent of the total number of samples, it is considered an anomaly. Finally, the lists of detected anomalies from each window are combined to get the final prediction using a similar criterion: an index is considered an anomaly if it appears in at least $\beta$ percent of the total number of overlapping windows, which are estimated by dividing the window size by step size. Here, $\alpha$ and $\beta$ are hyperparameters, which can be tuned to improve performance.
@@ 4.2 Detector: Finding Anomalies through Forecasting
As depicted in Figure 2, the first step of a typical ML pipeline is to train an ML model on a collection of time series. From Gruver et al. (2023), pretrained LLMs are capable of forecasting time series, allowing us to jump straight to the inference phase.
@@
Post-processing. For each time point $t$, we now have multiple forecasted values $\hat{x}_{t}$ in different windows when the horizon is larger than $1$, concretely $\{(\hat{x}^{i+h}_{t})\}$. We take the median from the collection as the final predicted value for $\hat{x}_{t}$. Furthermore, to increase the reliability of the prediction, we take $n$ samples from the distribution for each window $i$. Therefore, each $x_{t}$ has $n$ samples. To map it back to a univariate time series, we explore the results by taking the mean, median, $5^{th}$-percentile, and $95^{th}$-percentile as values. For the purpose of anomaly detection, an extreme forecast value could indicate a precursor to an anomaly; therefore acute values can be informative (see Section 3). Now, we have reconstructed the time series as $\mathbf{\hat{X}}=(\hat{x}_{1},\hat{x}_{2},\dots,\hat{x}_{T})$.
We proceed to compute the discrepancy between $\mathbf{X}$ and $\mathbf{\hat{X}}$. A large deviation indicates the presence of an anomaly. We denote this discrepancy as an error signal $e$, through computation of point-wise residuals, given their simplicity and ease of interpretation. We explore the usage of absolute difference suggested by Hundman et al. (2018) $e_{t}=\left|x_{t}-\hat{x}_{t}\right|$. Moreover, we explore how other functions, such as squared difference $e_{t}=(x_{t}-\hat{x}_{t})^{2}$ will help reveal the location of anomalies. More complex functions that capture the difference between two signals, such as dynamic time warping Müller (2007) can be used. However, Geiger et al. (2020) shows that discrepancies found with point-wise errors are sufficient for this purpose. Moreover, we apply an exponentially weighted moving average to reduce the sensitivity of the detection algorithm Hundman et al. (2018). Error values that surpass the threshold are considered anomalous. We use a sliding window approach to compute the threshold to help reveal contextual anomalies that are abnormal compared to the local neighborhood. As such, we assign the window size and step size to $T/3$ and $T/10$ respectively. We set a static threshold for each sliding window as four standard deviations away from the mean. These hyperparameters were chosen based on preliminary empirical results that agree with previous settings in other approaches Hundman et al. (2018); Geiger et al. (2020); Wong et al. (2022).
@@ 5 Evaluation
Datasets. We examined SigLLM on 11 datasets with known ground truth anomalies. These datasets were gathered from a wide range of sources, including a satellite telemetry signal corpus from NASA  [^6] https://github.com/khundman/telemanom that includes two sub-datasets: SMAP and MSL; Yahoo S5  [^7] https://webscope.sandbox.yahoo.com/catalog.php?datatype=s&did=70 which contains four sub-datasets: A1, which is based on real production traffic to Yahoo systems, and three others (A2, A3, and A4) which have been synthetically generated; and NAB  [^8] https://github.com/numenta/NAB which includes multiple types of time series data from various application domains. We consider five sub-datasets: Art, AWS, AdEx, Traf, and Tweets. In total, these datasets contain 492 univariate time series and 2,349 anomalies. The properties of each dataset, including the number of signals and anomalies, the average signal length, and the average length of anomalies are presented in Table 2. The table makes clear how properties differ between datasets; for instance, the NASA and NAB datasets contain anomalies that are longer than those in Yahoo S5, and the majority of anomalies in Yahoo S5’s A3 & A4 datasets are point anomalies.
@@
These models span a wide range of underlying detection methods, which increases successful detection coverage.
Metrics. We utilized anomaly detection-specific metrics for time series data Tatbul et al. (2018); Alnegheimish et al. (2022). Namely, we looked at the F1-Score, under which both partial and full anomaly detection are considered correct identification.
Hyperparameters. For Prompter, because of GPU capacity, the maximum input window length of SMAP and MSL is 500 values (for other datasets, it is 200 values). We chose a step size such that, on average, a value was contained in 5 overlapping windows (i.e, 100 steps for SMAP and MSL, and 40 for others). For Detector, we set the window size to 140 and the step size to 1. With a rolling window strategy of step size 1, it is important to keep the windows as small as possible while ensuring it is large enough to make useful predictions, as more context tends to be useful for LLMs. Our preliminary results suggested 140 was as performative as 200, and better than 100. We set the horizon to 5.
Computation. For GPT, we used GPT-3.5-turbo due to its superior performance on time series data (demonstrated by Gruver et al. (2023)) and its affordability. For Mistral, we used the publicly available model hosted by HuggingFace  [^9] https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.2 on an Intel i9-7920X 24 CPU core processor and 128GB RAM machine with 2 dedicated NVIDIA Titan RTX 24GB GPUs. To compute the benchmark. we set up an MIT Supercloud instance, with an Intel Xeon processor of 10 CPU cores (9 GB RAM per core) and one NVIDIA Volta V100 GPU with 32 GB memory.
@@ 5.1 Are large language models effective anomaly detectors for univariate time series?
After running the models on the entire datasets, we computed the precision, recall, and F1 scores, shown in Table 3. Overall, Mistral showcased better results than GPT for the Prompter method, with a 2×2$\times$2 × improvement in F1 score. In addition, Detector performed better overall than Prompter. We investigate each approach below.
@@
Detector Since we sampled multiple instances from the probability distribution, namely 10 samples in our experiments, we obtained a possible range of values that could be assigned to each particular point in time. We studied a multiple aggregation function to recreate a one-dimensional signal. Table 4 shows the detection F1 score when the signal is recreated from mean, median, $5^{th}$-percentile, and $95^{th}$-percentile values of the predicted distribution. Moreover, we consider different error scores under smoothing operation and without.
We can see that on average, squared error with smoothing on the median signal produced the best score, even though it was not the best performer on any individual dataset. The best configuration proved to be different for almost every dataset, with squared error producing the best results on Yahoo S5, and absolute error on NAB. One interesting observation is that the signal reconstructed from the $5^{th}/95^{th}$-percentile showcased higher potential in revealing the location of anomalies.
@@
Figure 7: Examples of successfully identified anomalies by Detector. Even though the model did not capture the trend present in synthetic_58, it still managed to find the anomalous intervals.
@@ 5.2 How does the SigLLM compare to existing approaches?
LLM-based methods can outperform transformer-based methods by 12.5%. Comparing Detector to AT, which is a transformer-based method, we see Detector outperforms AT in 7 out of the 11 datasets. However, this observation is strictly noticeable for Detector.
LLM-based methods can perform surprisingly well reaching an F1 score higher than MAvg by 14.6% and is only 10.9% away from ARIMA which is a reasonable model. Moreover, if we consider Detector only, the model surpasses MAvg performance in 6 out of 11 datasets, and ARIMA in 4 out of 11 dataset. We can best see the potential of Detector in the Tweets dataset, where the gap between the LLM’s result and the highest result (MAvg) is minimal, reaching only $1.8\%$.
AER, the best deep learning model is 30% better than LLM-based approaches. Deep learning methods still perform better than LLM-based methods by 18% on average. Observing the results for Detector, it is still lacking in comparison to deep learning models. With a gap as big as 30% to the highest performing model (AER) and as little as 5% when comparing to the least performative model (VAE). Moreover, LLMs do not hold the highest score for any dataset. It is clear that there is a significant gap here, and an opportunity for improvement.
In the following section, we aim to address where the LLMs succeeded and where they fell short.
@@ 5.3 What are the success and failure cases and why?
Figures 7 and 7 illustrate example outputs for both the Prompter and Detector methods, respectively. We focus on Mistral since it yielded better overall results than GPT as depicted in Table 3.
Success Cases. For both signals shown in Figure 7, Detector correctly identified all the anomalies, with only one false alarm in the second signal. The Prompter approach is able to detect outliers and extreme values in its local context a lot more effectively than anomalies that are buried within the signal. For example, if the window is “244 , 309 , 2 , 462”, both GPT and Mistral point to “2” as the observed anomaly. However, this problem becomes more ambiguous when the context is larger. As seen in Figure 7 on Twitter_volume_AMZN signal, Prompter identified some local extreme values as anomalous. Oddly enough, it also missed some anomalies that were global outliers.
Failure Cases. Even though Detector correctly identified all anomalies in the example shown in Figure 7, the forecast itself struggled to capture the non-stationary aspects of the signal, particularly its trend. This is due to the sensitivity of LLMs to the context length. A window size larger than 140 is needed to capture this property. While it did not impact the detection in this case, it might be the reason behind failure cases in other signals.
With Prompter, the number of false alarms raised was significantly high, with an average precision of 0.219. Using the filtering method described in Section 4.1 does not eliminate the false positives. An alternative could be to use log probabilities as a measure of confidence for filtering. We recommend exploring this avenue in future work.
@@ 6 Discussion
Prompting challenges Over a three-month experimental period, various prompts were employed, as laid out in Table 1. It is evident that both GPT and Mistral fail to produce the desired responses unless a chat template is applied, attributing roles to the user and the system. Furthermore, to ensure the exclusivity of numerical values in the generated responses, beyond merely specifying in the prompt to “just return numbers”, we adjusted the likelihood of non-numerical tokens appearing in the output generated by the LLMs.
@@
Under the ‘find indices’ prompt, GPT may generate lists of indices; however, these indices frequently surpass the sequence length. Conversely, Mistral yields values instead of indices when utilizing the same prompt. Therefore, we alter the prompt to include ‘find values’ rather than ‘find indices’ for our experiment.
@@
Unlike Mistral, GPT outputs “repetitive prompt” error when presented with a series of identical values within a window, which was particularly evident in NASA datasets (there are 23 signals in SMAP and 13 in MSL with this error, affecting up to 85% of the windows). In this experiment, we deemed such windows as having no detectable anomalies, and hence, obtain a true positive of zero.
@@
Addressing Memorization. Large language models are trained on a vast amount of data. The training data for most models – for instance those provided by OpenAI – is completely unknown to the general public, which makes evaluating these models nuanced. Given that large language models, especially GPT models Chang et al. (2023), are notorious for memorizing training data Biderman et al. (2023), how do we ensure that there was not any data and label leakage of the benchmark datasets used? We posit that our transformation of the time series data into its string representation is unique, essentially making the input time series different from its original form and reducing the chances of blatant memorization. Moreover, unlike with the forecasting task, the task of anomaly detection is not inherent to the training convention used, which is next token prediction.
@@
Figure 8: Recorded time for Prompter and Detector. (left) On average, Detector takes the longest to infer, almost double the time of Prompter. (right) Distribution of signal length and how long it takes to execute.
@@
Practicality of Usage. The appeal of using LLMs for this task resides in their ability to be used in zero-shot, without necessitating any fine-tuning. However, this property is bottlenecked by the latency time. Figure 8 illustrates the average time it takes to use LLMs for each of the suggested approaches. It is unreasonable to wait half an hour to two hours for the model to produce a response, especially when deep learning models take less than an hour to train. Since these experiments were ran in an offline setting, we can expect that real time deployment would be more sensible, since the context size is much smaller, eliminating the need for rolling windows. In other cases, a single window would be sufficient, which takes approximately 5 seconds to infer depending on the window size.
@@
In total, running Prompter experiments using the gpt-3.5-turbo-instruct version has cost us approximately \$834.11, which is an average of \$1.69 per signal. As for Detector, we ran a small-scale experiment where we sampled 22 signals of the data, roughly 5%. The total reached \$95.08, which is an average of \$4.3 per signal. Detector is a more expensive method to run compared to Prompter.

### v3

@@ 4 SigLLM: Detecting Anomalies in Signals using Large Language Models
4.1 Prompter: Finding Anomalies through Prompting
@@
Under this prompt, the LLM generates a list of values it delineates as point-wise anomalies. It is noteworthy that the GPT-3.5-turbo model is capable of directly outputting anomalous indices using the prompt presented in Table 1, while Mistral lacks this ability, as shown in trial #5. To maintain consistency across our experiments, we conducted experiments on both models using the same prompt mentioned above.
As explained in Section III.A, we adopt the rolling windows method, segmenting the time series into rolling windows before inputting it into the LLMs. For each window, we generate 10 samples from the output probability distribution. For each sample containing values deemed anomalous by LLMs, we collect all indices of the window corresponding to those values. Then, the 10 lists of indices are merged together: if an index appears in at least $\alpha$ percent of the total number of samples, it is considered an anomaly. Finally, the lists of detected anomalies from each window are combined to get the final prediction using a similar criterion: an index is considered an anomaly if it appears in at least $\beta$ percent of the total number of overlapping windows, which are estimated by dividing the window size by the step size. Here, $\alpha$ and $\beta$ are hyperparameters, which can be tuned to improve performance.
@@ 4.2 Detector: Finding Anomalies through Forecasting
As depicted in Figure 2, the first step in a typical ML pipeline involves training an ML model on a collection of time series. From Gruver et al. (2023), pretrained LLMs are capable of forecasting time series, allowing us to jump straight to the inference phase.
@@
Post-processing. For each time point $t$, we now have multiple forecasted values $\hat{x}_{t}$ in different windows when the horizon is larger than $1$, concretely $\{(\hat{x}^{i+h}_{t})\}$. We take the median from the collection as the final predicted value for $\hat{x}_{t}$. Furthermore, to increase the reliability of the prediction, we take $n$ samples from the distribution for each window $i$. Therefore, each $x_{t}$ has $n$ samples. To map this back to a univariate time series, we explore the results by taking the mean, median, $5^{th}$-percentile, and $95^{th}$-percentile as values. For the purpose of anomaly detection, an extreme forecast value could indicate a precursor to an anomaly; therefore, acute values can be informative (see Section 3). Now, we have reconstructed the time series as $\mathbf{\hat{X}}=(\hat{x}_{1},\hat{x}_{2},\dots,\hat{x}_{T})$.
We next compute the discrepancy between $\mathbf{X}$ and $\mathbf{\hat{X}}$. A large discrepancy indicates the presence of an anomaly. We denote this discrepancy as an error signal $e$ by computing point-wise residuals, given their simplicity and ease of interpretation. We explore the usage of absolute difference suggested by Hundman et al. (2018) $e_{t}=\left|x_{t}-\hat{x}_{t}\right|$. Moreover, we explore how other functions, such as squared difference $e_{t}=(x_{t}-\hat{x}_{t})^{2}$ will help reveal the location of anomalies. More complex functions that capture the difference between two signals, such as dynamic time warping Müller (2007) can be used. However, Geiger et al. (2020) shows that discrepancies found with point-wise errors are sufficient for this purpose. Moreover, we apply an exponentially weighted moving average to reduce the sensitivity of the detection algorithm Hundman et al. (2018). Error values that surpass the threshold are considered anomalous. We use a sliding window approach to compute the threshold to help reveal contextual anomalies that are abnormal compared to the local neighborhood. As such, we assign the window size and step size to $T/3$ and $T/10$ respectively. We set a static threshold for each sliding window as four standard deviations away from the mean. These hyperparameters were chosen based on preliminary empirical results that agree with previous settings in other approaches Hundman et al. (2018); Geiger et al. (2020); Wong et al. (2022).
@@ 5 Evaluation
Datasets. We examined SigLLM on 11 datasets with known ground truth anomalies. These datasets were gathered from a wide range of sources, including a satellite telemetry signal corpus from NASA  [^6] https://github.com/khundman/telemanom that includes two sub-datasets: SMAP and MSL; Yahoo S5  [^7] https://webscope.sandbox.yahoo.com/catalog.php?datatype=s&did=70, which contains four sub-datasets: A1, which is based on real production traffic to Yahoo systems, and three others (A2, A3, and A4) which have been synthetically generated; and NAB  [^8] https://github.com/numenta/NAB, which includes multiple types of time series data from various application domains. We consider five sub-datasets: Art, AWS, AdEx, Traf, and Tweets. In total, these datasets contain 492 univariate time series and 2,349 anomalies. The properties of each dataset, including the number of signals and anomalies, the average signal length, and the average length of anomalies, are presented in Table 2. The table makes clear how properties differ between datasets; for instance, the NASA and NAB datasets contain anomalies that are longer than those in Yahoo S5, and the majority of anomalies in Yahoo S5’s A3 & A4 datasets are point anomalies.
@@
These models use a wide range of underlying detection methods, which increases our anomaly detection coverage overall.
Metrics. We utilized anomaly detection-specific metrics for time series data Tatbul et al. (2018); Alnegheimish et al. (2022). Namely, we looked at the F1 score, under which both partial and full anomaly detection are considered correct identification.
Hyperparameters. For Prompter, GPU capacity means that the maximum input window length of SMAP and MSL is 500 values (for other datasets, it is 200 values). We chose a step size such that, on average, a value was contained in 5 overlapping windows (i.e, 100 steps for SMAP and MSL, and 40 for others). For Detector, we set the window size to 140 and the step size to 1. With a rolling window strategy of step size 1, it is important to keep the windows as small as possible while still ensuring that they are large enough to make useful predictions, as more context tends to be useful for LLMs. Our preliminary results suggested that a window size of 140 was as performative as a window size of 200, and was better than a window size of 100. We set the horizon to 5.
Computation. For GPT, we used GPT-3.5-turbo due to its superior performance on time series data (demonstrated by Gruver et al. (2023)) and its affordability. For Mistral, we used the publicly available model hosted by HuggingFace  [^9] https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.2 on an Intel i9-7920X 24 CPU core processor and 128GB RAM machine with 2 dedicated NVIDIA Titan RTX 24GB GPUs. For benchmarking, we use Intel Xeon processor of 10 CPU cores (9 GB RAM per core) and one NVIDIA Volta V100 GPU with 32 GB memory.
@@ 5.1 Are large language models effective anomaly detectors for univariate time series?
After running the models on the full datasets, we computed the precision, recall, and F1 scores, shown in Table 3. Overall, Mistral achieved better results than GPT for the Prompter method, with a 2×2$\times$2 × improvement in F1 score. In addition, Detector performed better overall than Prompter. We investigate each approach below.
@@
Detector Since we sampled multiple instances from the probability distribution (namely 10 samples in our experiments), we obtained a possible range of values that could be assigned to each particular point in time. We studied a multiple aggregation function to recreate a one-dimensional signal. Table 4 shows the detection F1 score when the signal is recreated from mean, median, $5^{th}$-percentile, and $95^{th}$-percentile values of the predicted distribution. Moreover, we consider different error scores under smoothing operation and without.
We can see that, on average, squared error with smoothing on the median signal produced the best score, even though it was not the best performer on any individual dataset. The best configuration proved to be different for almost every dataset, with squared error producing the best results on Yahoo S5, and absolute error on NAB. One interesting observation is that the signal reconstructed from the $5^{th}/95^{th}$-percentile showcased higher potential in revealing the location of anomalies.
@@
Figure 7: Examples of anomalies successfully identified by Detector. Even though the model did not capture the trend present in synthetic_58, it still managed to find the anomalous intervals.
@@ 5.2 How does the SigLLM compare to existing approaches?
LLM-based methods can outperform transformer-based methods by 12.5%. Comparing Detector to AT, which is a transformer-based method, we see Detector outperforms AT in 7 out of the 11 datasets. However, this observation is only relevant to Detector.
LLM-based methods can perform surprisingly well. Our methods achieved an F1 score 14.6% higher than that of MAvg, and only 10.9% lower than that of ARIMA which is a reasonable model. Moreover, the Detector pipeline alone surpasses MAvg performance in 6 out of 11 datasets, and ARIMA in 4 out of 11 dataset. We can best see the potential of Detector in the Tweets dataset, where the gap between the LLM’s result and the highest result (from MAvg) is minimal, at only $1.8\%$.
AER, the current best deep learning model, is 30% better than LLM-based approaches. Deep learning methods still perform better than LLM-based methods by 18% on average. Looking more closely at Detector shows that it does not perform as well as deep learning models, achieving a 30% lower score than that of the highest performing model (AER) and a 5% lower score than the that of least performative model (VAE). Moreover, LLMs do not hold the highest score for any dataset. It is clear that there is a significant gap here, and an opportunity for improvement.
In the following section, we aim to address where the LLMs succeeded at this problem and where they fell short.
@@ 5.3 What are the success and failure cases and why?
Figures 7 and 7 illustrate example outputs for both the Prompter and Detector methods, respectively. We focus on Mistral since it yielded better overall results than GPT, as depicted in Table 3.
Success Cases. For both signals shown in Figure 7, Detector correctly identified all the anomalies, with only one false alarm in the second signal. The Prompter approach is able to detect outliers and locally extreme values a lot more effectively than anomalies that are buried within the signal. For example, if the window is “244 , 309 , 2 , 462”, both GPT and Mistral point to “2” as the observed anomaly. However, this problem becomes more ambiguous when the context is larger. As seen in Figure 7 on Twitter_volume_AMZN signal, Prompter identified some locally extreme values as anomalous. Oddly enough, it also missed some anomalies that were global outliers.
Failure Cases. Even though Detector correctly identified all anomalies in the example shown in Figure 7, the forecast itself struggled to capture the non-stationary aspects of the signal, particularly its trend. This is due to the sensitivity of LLMs to context length. A window size larger than 140 is needed to capture this property. While it did not impact the detection in this case, this may explain failure cases in other signals.
Prompter raised a large number of false alarms, with an average precision of 0.219. Using the filtering method described in Section 4.1 does not eliminate false positives. An alternative strategy could be to use log probabilities as a measure of confidence for filtering. We recommend exploring this avenue in future work.
@@ 6 Discussion
Prompting Challenges. Over a three-month experimental period, various prompts were employed, as laid out in Table 1. It is evident that both GPT and Mistral fail to produce the desired responses unless a chat template is applied that attributes roles to the user and the system. Furthermore, to ensure the exclusivity of numerical values in the generated responses, in addition to specifying in the prompt to “just return numbers,” we adjusted the likelihood of non-numerical tokens appearing in the output generated by the LLMs.
@@
Under the ‘find indices’ prompt, GPT may generate lists of indices; however, these indices frequently surpass the sequence length. Conversely, Mistral yields values instead of indices when utilizing the same prompt. Therefore, for our experiment, we altered the prompt to include ”find values” rather than ”find indices.”
@@
Unlike Mistral, GPT outputs a “repetitive prompt” error when presented with a series of identical values within a window. This happened particularly for NASA datasets (there are 23 signals in SMAP and 13 in MSL with this error, affecting up to 85% of the windows). In this experiment, we deemed such windows as having no detectable anomalies, obtaining a true positive of zero.
@@
Addressing Memorization. Large language models are trained on a vast amount of data. The training data for most models – for instance, those provided by OpenAI – is completely unknown to the general public, which makes evaluating these models a nuanced problem. Given that large language models, especially GPT models Chang et al. , are notorious for memorizing training data Biderman et al. (2023), how do we ensure that there was no data and label leakage for the benchmark datasets used? We posit that our transformation of the time series data into its string representation is unique, essentially making the input time series different from its original form and reducing the chances of blatant memorization. Moreover, unlike with the forecasting task, the task of anomaly detection is not inherent to the training convention used, which is next token prediction.
@@
Figure 8: Recorded time for Prompter and Detector. (left) On average, Detector takes the longest to infer, almost double the time of Prompter. (right) Distribution of signal length and execution time.
@@
Practicality of Usage. The appeal of using LLMs for this task lies in their ability to be used in zero-shot, without necessitating any fine-tuning. However, this property is bottlenecked by the latency time. Figure 8 illustrates the average time it takes to use LLMs for each of the suggested approaches. It is unreasonable to wait half an hour to two hours for the model to produce a response, especially when deep learning models take less than an hour to train. Since these experiments were run in an offline setting, we can expect that real-time deployment would be more sensible, especially since the context size is much smaller, eliminating the need for rolling windows. In other cases, a single window would be sufficient, which takes approximately 5 seconds to infer depending on the window size.
@@
In total, running Prompter experiments using the gpt-3.5-turbo-instruct version has cost us approximately \$834.11 – an average of \$1.69 per signal. For Detector, we ran a small-scale experiment where we sampled 22 signals of the data, roughly 5%. The total reached \$95.08 – an average of \$4.3 per signal – making Detector a more expensive method than Prompter.